A pension fund owns 12,000 shares in a stock mutual fund and a mutual bond fund. Currently, the stock fund sells for $12 per share and the bond fund sells for $15 per share. If the value of the securities is $165,000, how many shares of each does the pension fund own?
Answer by checkley71(8403)

15X+12(12,000-X)=165000 OR 15X+144000-12X=165000 OR 3X=165000-144000 OR
3X=21000 OR X=21000/3 OR X=7000 SHARES @ $15 & 12000-7000=5000 @ $12
PROOF 7000*15+5000*12=165000 OR 105000+60000=165000 OR 165000=165000